OpenStudy (anonymous):

The three sides of a right triangle are x, x+3, and 2x-3.(

OpenStudy (jdoe0001):

ok, now we know what the three sides are I wouldn't be losing sleep over that now

OpenStudy (anonymous):

well i need to find x

OpenStudy (anonymous):

read the whole question @jdoe0001 ^^^^

OpenStudy (jdoe0001):

hmm that was added just now

OpenStudy (anonymous):

well anyways, what is x?

OpenStudy (anonymous):

#phythagorean theorem

OpenStudy (jdoe0001):

 \(\bf \textit{pythagorean}=c^2=a^2+b^2\implies c=\pm \sqrt{a^2+b^2} \\ \quad \\ 2x-1=\pm \sqrt{(x)^2+(x+3)^2}\) solve for "x"

OpenStudy (anonymous):

how do i solve for x?

OpenStudy (jdoe0001):

expand the binomial on the left-side, and the one on the right-side then simplify

OpenStudy (anonymous):

What to you mean by expand? Do the multiplication so it's 4x+1?

OpenStudy (jdoe0001):

pythagorean theorem I meant anyhow you'd use the pythagorean theorem, since it's a right-triangle

OpenStudy (jdoe0001):

hmmm

OpenStudy (jdoe0001):

shoot I got those vlaues a bit off... .lemme redo that

OpenStudy (anonymous):

would it be 4x+1?

OpenStudy (jdoe0001):

 \(\bf \large { \textit{pythagorean theorem}=c^2=a^2+b^2 \\ \quad \\ (2x-3)^2=(x)^2+(x+3)^2 }\)

OpenStudy (anonymous):

4x+1=x^2+2x+9?

OpenStudy (jdoe0001):

hmmm actually you'd need to use FOIL or just multply the binomials bear in mind that \(\bf (2x-3)^2=(x)^2+(x+3)^2\implies (2x-3)(2x-3)=(x)^2+(x+3)(x+3)\)

OpenStudy (jdoe0001):

hmm got a bit truncated.... ok so \(\bf (2x-3)^2=(x)^2+(x+3)^2 \\ \quad \\ \implies (2x-3)(2x-3)=(x)^2+(x+3)(x+3) \)

OpenStudy (anonymous):

ok so i got 4x^2-12x+9=x^4+6x+9

OpenStudy (anonymous):

is that correct?

OpenStudy (anonymous):

@jdoe0001

OpenStudy (jdoe0001):

hmm close one sec

OpenStudy (jdoe0001):

hmmm wait a sec.. shoot... lemme fix that quick

OpenStudy (jdoe0001):

\(\bf (2x-3)^2=(x)^2+(x+3)^2 \\ \quad \\ (2x-3)(2x-3)=(x)^2+(x+3)(x+3) \\ \quad \\ 4x^2-12x+9=x^2+x^2+6x+9 \\ \quad \\ 4x^2-12x+9=2x^2+6x+9 \\ \quad \\ 2x^2-18x=0\implies 2x(x-9)=0\impliedby common\ factoring \\ \quad \\ 2x(x-9)=0\implies \begin{cases} 2x=0\to x=\frac{0}{2}\to &x=0\\ x-9=0\to &x=9 \end{cases}\)

OpenStudy (anonymous):

so x is 9, so the sides are 9, 12, and 15?

OpenStudy (jdoe0001):

yeap

OpenStudy (anonymous):

Ok thanks

OpenStudy (jdoe0001):

yw